OpenStudy (anonymous):

What is the first step in solving this system of equations by elimination? 2x – 2y = 5 5x + 2y = 9

OpenStudy (anonymous):

eliminating the y's

OpenStudy (anonymous):

Is this multiple choice or written response?

OpenStudy (anonymous):

Multiple choice

OpenStudy (anonymous):

A. Subtract the two equations. B. Add the two equations. C. Solve for y. D. Multiply each equation by 2.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

dude I will follow you and give you a medal if you help me

OpenStudy (anonymous):

Ok thanks! Can you help me on 2 more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks! Which of the following would be the best first step to solve the system of equations? 2x – 9y = 6 4x – 3y = –7

OpenStudy (anonymous):

A. Multiply the first equation by 4. B. Multiply the second equation by 2. C. Multiply the first equation by –3. D. Multiply the second equation by –3.

OpenStudy (anonymous):

multiply the second equation by -3

OpenStudy (anonymous):

you want to get x or y the same number but opposite signs first so those can be eliminated

OpenStudy (anonymous):

OK one more one minute I have to type the equation

OpenStudy (anonymous):

k

OpenStudy (anonymous):

How many solutions will the following system of linear equations have? -2x+4y=-7 \[y=-\frac{ 1 }{ 2 }x+\]

OpenStudy (anonymous):

it was +5

OpenStudy (anonymous):

A. 0 B. 1 C. 2 D. infinite

OpenStudy (anonymous):

Nvm I submitted it and got a 90. Thanks dude! you are the best

OpenStudy (anonymous):

sorry was gonna get someone to help you >.< my guess was 2 but wasnt completly sure